OpenStudy (anonymous):

in the quadratic equation 25^2-36=0, i thought that a=25^2, b=-36 and c=0.

OpenStudy (anonymous):

No the equation is: ax^2 +bx + c if it does not have one of these you can replace it with 0

OpenStudy (anonymous):

so for you it would be: 25^2 + 0 -36= 0

OpenStudy (anonymous):

"b" will always have the variable "x" next to it if it doesn't then it is "c" not "b"

OpenStudy (anonymous):

Do you know how to solve it?

OpenStudy (anonymous):

i am going to try I was just thrown off with the a, b, and c.

OpenStudy (anonymous):

can you solve itin detail

OpenStudy (anonymous):

No prob :)

OpenStudy (anonymous):

what do you mean 25^2 there suppose to be an x did u miss it?

OpenStudy (anonymous):

IS it 25x^2? or 2x^2 or 5x^2

OpenStudy (anonymous):

Well here is what you would do: since the equation is ax + bx + c you would multiply "c" by "a" ONLY if "a" is something besides 1 or -1, if it's negative one then you would multiply both sides by -1 to cancel it out. So after you make sure a is not something beside 1 then find all the factors of "c" for example if "c" was 20 you would do this: 2 * 10 1 * 20 4 * 5

OpenStudy (anonymous):

25x^2-36=0 is the problem

OpenStudy (anonymous):

Oh ok now I can teach you accurately :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So first of since "a" is not 1 or -1 we multiply it by "c"

OpenStudy (anonymous):

Giving us 900 now we would list all the factors of 900 and of all of those we would pick the one that equals 0

OpenStudy (anonymous):

It would be -900 btw

OpenStudy (anonymous):

900 what do you mean? I'm really confused now

OpenStudy (anonymous):

Cause you told me that "a" was 25 and we have to multiply 25 by -36

OpenStudy (anonymous):

Which would give me -900

OpenStudy (anonymous):

Unless its a typo that's how it is

OpenStudy (anonymous):

25x^2-36=0 I may have typed it incorrectly.

OpenStudy (anonymous):

I'll just tell u what u should do

OpenStudy (anonymous):

Well here is what you would do: since the equation is ax + bx + c you would multiply "c" by "a" ONLY if "a" is something besides 1 or -1, if it's negative one then you would multiply both sides by -1 to cancel it out. So after you make sure a is not something beside 1 then find all the factors of "c" for example if "c" was 20 you would do this: 2 * 10 1 * 20 4 * 5 And if "b" was 9 then you would pick the pair that would add up to 9 so you would pick 5 and 4. Now you would put both variable in parenthesis with an "x" next to them like so: (x + 4) and (x + 5) now you would make them equal 0 x+4=0 and x+5=0 Then do the inverse operation and subtract from both sides so: x= -4 and x= -5 So your parabola will intersect the x intercept at (-4, 0) and (-5, 0)

OpenStudy (anonymous):

the onlie math class returned the solution as ( -6/5, 6/5)

OpenStudy (anonymous):

You have flvs?

OpenStudy (anonymous):

what is fls?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

nvm?

OpenStudy (anonymous):

nevermind means nvm

OpenStudy (anonymous):

ok im not real hip

OpenStudy (anonymous):

what did you mean by flvs???